### v2

@@ 8.1 Finding a large planar part
We claim the following: If at least $1000g$ inner vertices of $W_{1}$ are connected to the outside, then $G$ has genus at least $g$. The strategy is similar to the argument in Kociumaka and Pilipczuk [22, Lemma 5.3]; we summarize the proof. If at least $1000g$ inner vertices of $W_{1}$ are connected to the outside, then a set $U$ of $g$ inner vertices of $W_{1}$ are connected to the outside, and at pairwise distance at least 16 in $W_{1}$. This implies that $G$ contains, as a minor, the graph $J_{g}$ obtained from $g$ copies of $K_{5}$ by subdividing an edge from each copy with a degree-two vertex and identifying these $g$ new vertices into a single vertex, the apex of $J_{g}$; see Figure 8.2. This graph has genus at least $g$, by the main result of an article by Miller [27, Theorem 1]. This proves the claim.
@@
Figure 8.2: Illustration of the proof of the claim in Proposition 8.1. The wall $W_{1}$ (left), and a schematic view of the $J_{p}$ minor (right), illustrated with $g=6$. Let $K_{5}^{-}$ be the graph $K_{5}$ with one edge removed. If $W_{1}$ has many vertices connected to the outside, then a set $U$ of at least $g$ vertices connected to the outside (represented as big disks on the left) are pairwise distant. In the neighborhood of these vertices, we build a $K_{5}^{-}$ minor for each vertex in $U$, in which the vertex of $U$ is the “central” vertex (the thick paths need to be contracted to obtain copies of $K_{5}^{-}$). Using the fact that the neighborhoods are far apart, we build a $J_{g}$-minor of $G$ (right), in which the apex results from the contraction of the subdivided wall $\dot{W}_{1}$ minus the union of the $K_{5}^{-}$ minors, and each path connecting $u\in U$ to $\dot{W}-\dot{W}_{1}$ (which is connected by assumption) is contracted, except its first edge, to connect $u$ to the apex.
@@ 8.2 Finding an irrelevant vertex
Let $\mathscr{C}$ be a 2-complex with $c\geq 1$ simplices. Let $G$ be a graph and $\gamma$ be a cycle in $G$ such that one connected component of $G-\gamma$ is planar and contains a subdivision of the $(200c\times 200c)$-wall. Let $v$ be the central vertex of this wall. Then $G$ is embeddable on $\mathscr{C}$ if and only if $G-v$ is.
The proof is inspired from an article by Kociumaka and Pilipczuk [22, Section 5.3], but we provide a slightly different argument, also because we need to handle a 2-complex, not a surface. We use more concepts of topology, in particular homotopy; see, e.g., Stillwell [38].
@@
On a surface of genus $g$, a set of disjoint, simple closed curves belong to at most $23g+$ 2 (free) homotopy classes (here we regard a closed curve and its reversal to be in the same homotopy class).
@@
We need some preparations. A circular wall $W$ [22, Figure 8] of height $h$ and circumference $\ell$ is a 3-regular graph that, in some embedding of $W$ in the plane, is represented as the union of $h$ vertex-disjoint cycles, called circles, organized in a concentric way, such that any two consecutive circles are connected by $\ell$ radial edges; in any two successive layers, the radial edges are interleaved.
@@
Let $H$ be the connected component of $G-\gamma$ that is planar and contains a subdivision of the $(200g\times 200g)$-wall. In the following, we denote by $\varphi$ the plane embedding of $H$ such that $v$ is the central vertex of this wall. Then $H$ contains a subdivision $\dot{W}$ of a circular wall $W$ of height $520c+$ 5 and circumference $12c+$ 1, so that $v$ is located inside the inner circle of this circular wall in $\varphi$; see Kociumaka and Pilipczuk [22, Figures 8 and 9].
@@
If $G$ is embeddable on $\mathscr{C}$, then obviously $G-v$ is also embeddable on $\mathscr{C}$; the hard part is the reverse direction. So let us consider an embedding $\psi$ of $G-v$ on $\mathscr{C}$. This induces embeddings of $\dot{W}$ and $W$ on $\mathscr{C}$. At most $c$ vertices or edges of $W$ are mapped, under $\psi$, to a singular vertex of $\mathscr{C}$, and the isolated edges of $\mathscr{C}$ intersect the images of at most $c$ edges of $W$. Thus, we obtain a subdivision $\dot{W}^{\prime}$ of a circular wall $W^{\prime}$ of height $518c+$ 5 and circumference one that does not intersect any singular vertex or isolated edge of $\mathscr{C}$.
@@
Now, let $\mathscr{S}$ be the detached surface of $\mathscr{C}$, and let $\hat{\mathscr{S}}$ be the surface obtained by attaching a handle to every boundary component of $\mathscr{S}$ and to each point of $\mathscr{S}$ not on its boundary that corresponds, in $\mathscr{C}$, to a singular vertex. At most $2c$ handles were attached, so $\hat{\mathscr{S}}$ has a natural cellular graph embedding with at most $3c$ edges, and thus has genus at most $3c$. Moreover, $\psi$ naturally corresponds to an embedding of $\dot{W}^{\prime}$ and $W^{\prime}$ on $\hat{\mathscr{S}}$, which we still denote by $\psi$.
By Lemma 8.4, since $\hat{\mathscr{S}}$ has genus at most $3c$, and $W^{\prime}$ has height $12(3*3c+2)+$ 1, there are three pairwise distinct circles of the subdivision $\dot{W}^{\prime}$ of the circular wall $W^{\prime}$ that are homotopic under $\psi$. They are two-sided, because any two one-sided homotopic closed curves cross. Let $C^{i}$, $C$, and $C^{o}$ be these three circles, in this order, starting from the innermost circle in $\varphi$ (thus $C^{i}$ is inside, and $C^{o}$ outside).
A bridge (of $\dot{W}$) is either (1) a connected component of $G-\{v\}-\dot{W}$, together with the edges joining it to $\dot{W}$, or (2) an edge of $G-\{v\}-\dot{W}$ whose endpoints are on $\dot{W}$. For a bridge of $\dot{W}$, the vertices of $\dot{W}$ that are endpoints of edges in the bridge are the attachment points of the bridge. We partition the vertices and edges of $G-v$ as follows: (1) the cycle $C$; (2) the part $I$ of $G-v$ that lies inside $C$ in $\varphi$; (3) the union $B$ of the bridges (of $\dot{W}$) that are mapped outside $C$ by $\varphi$, have at least one attachment point, but all of them lie on $C$; (4) the part $O$ of $G-v$ that lies in the same connected component of $G-v$ as $C$, but is disjoint from $C$, $I$, and $B$; (5) the union $R$ of the other connected components of $G-v$. We remark that $O$ is connected and contains $C^{o}$, because the circumference of $W$ is at least one.
We now claim that in $\psi$, the edges of $O$ that have an endpoint on $C$ leave $C$ all on the same side of $C$. For this, we distinguish two cases.
  • Assume that $\psi(C)$ is contractible on $\mathscr{S}$; it bounds a disk in $\mathscr{S}$, and thus also in $\mathscr{C}$. Since $O$ is connected, it is mapped, by $\psi$, either completely inside that disk, or completely outside, hence the claim holds.
@@
  • Otherwise, the three circles $\psi(C^{i})$, $\psi(C)$, and $\psi(C^{o})$ are homotopic but non-contractible. So any two of these closed curves bound an annulus in $\mathscr{S}$ [17, Lemma 2.4], and thus also in $\mathscr{C}$. There are two subcases:
@@
    • Assume that, under $\psi$, the annulus bounded by $C$ and $C^{i}$ does not contain $C^{o}$. Since $O$ is connected, contains $C^{o}$, and is disjoint from $C$ and $C^{i}$, it cannot enter the annulus bounded by $C$ and $C^{i}$. So indeed the claim holds.
    • Otherwise, under $\psi$, the circle $C^{o}$ lies in an annulus bounded by $C$ and $C^{i}$. Since $O$ is connected, contains $C^{o}$, and is disjoint from $C$ and $C^{i}$, it must lie in the annulus bounded by $C$ and $C^{o}$. So the claim also holds.
We consider the restriction $\psi^{\prime}$ of $\psi$ to $C\cup O\cup R$. We extend $\psi^{\prime}$ to an embedding $\psi^{\prime\prime}$ of $C\cup O\cup R\cup B$, by embedding $B$ on the same side of $C$ as $O$, close to $C$, as in $\varphi$. In $\psi^{\prime\prime}$, all the edges incident to $C$ leave $C$ all on the same side of $C$. Finally, we extend $\psi^{\prime\prime}$ to all of $G$, by embedding the remaining part, $I\cup\{v\}$, in a small neighborhood of $C$, on the side of $C$ not used by $O$, in the same way as $C\cup I\cup\{v\}$ is planarly embedded in $\varphi$. ∎
@@
    a rooted branch decomposition of $G$ of width $O(g)^{O(1)}$;
@@
In the first two cases, we are done. In the third case, by applying Proposition 8.3, we obtain a vertex $v$ such that $G$ embeds on $\mathscr{C}$ if and only if $G-v$ does. By iterating the same procedure a number of times that is at most the number of vertices of $G$, we necessarily reach case (1) or (2), which concludes. ∎
@@ 9 Computing an embedding
  • the images of the vertices of $H$ cover the singular points of $\mathscr{C}$;
  • the restriction of $H$ to the detached surface $\mathscr{S}$ is specified by (1) a cellular embedding of a graph $H^{\prime}$, represented by its (standard) combinatorial map, such that $E(H^{\prime})\subseteq E(H)$, (2) a map from $V(H^{\prime})$ to $V(H)$ describing how the vertices of $H^{\prime}$ are identified when attaching the points of $\mathscr{S}$ to recover the singular points of $\mathscr{C}$, and (3) a map from the images of the singular points of $\mathscr{C}$ on $\mathscr{S}$ to $V(H^{\prime})$, to specify which vertex of $H^{\prime}$ occupies that point. Moreover, we have $|V(H^{\prime})|\leq|V(H)|+c$;
  • each isolated edge $e$ of $\mathscr{C}$ contains vertices of $H$, edges of $H$, and “holes” (parts of $e$ not in the image of $H$), which we can represent by listing the corresponding sequence in order along $e$.
We remark that it is not a severe restriction to assume that $\mathscr{C}$ has no 3-book, since any graph embeds on a 3-book (Proposition 3.1) and no isolated vertex, since only isolated vertices of $G$ can be embedded on such isolated vertices. Also, we remark that the graph $H$ is needed because we may need to add edges to make the embedding cellular on the detached surface, and we may need to subdivide edges in order to avoid any edge to go through a singular vertex.
@@
    Since $G$ embeds on $\mathscr{C}$, for some complex $\mathscr{C}_{i}$, we have computed a sparse bounding graph for $G^{\prime}$ on $\mathscr{C}_{i}$ at the root of the rooted branch decomposition. We can use backtracking: at each node $\alpha$ of the rooted branch decomposition, we store the embedding of the union of $G_{\alpha}$ and of its bounding graph, which is a proper embedding. Eventually, the embedding of $G^{\prime}$ at the root of the recursion is not necessarily proper, but it becomes proper if we add isolated vertices mapped to singular points that are not covered by the embedding of $G^{\prime}$. Thus, from $(G^{\prime},\mathscr{C}_{i})$, we can return a proper embedding of the graph $G^{\prime}$, augmented with at most $c$ isolated vertices, into $\mathscr{C}_{i}$.
@@
In Theorem 1.1, if $\mathscr{C}$ has no 3-book and no isolated vertex, $G$ has no isolatex vertex and at most one connected component that is a path, then one can compute a proper embedding of a $5c$-homeomorph of $G$, augmented with at most $c$ isolated vertices, into $\mathscr{C}$ (using the data structure of Section 3.4), without overhead in the asymptotic running time.
@@
The decision algorithm iteratively removes vertices to $G$, getting a smaller equivalent instance. For the algorithm that actually computes an embedding, we now need recursion. The base case is given by Lemma 9.2. We assume that for some vertex $v$ of $G$, we have a proper embedding of a $5c$-homeomorph of $G-v$, augmented with at most $c$ isolated vertices, into $\mathscr{C}$. We need to recover a proper embedding of a $5c$-homeomorph of $G$, augmented with at most $c$ isolated vertices, into $\mathscr{C}$. We note that $v$ is the central vertex of a $(200c\times 200c)$-wall contained in one planar connected component of $G-\gamma$, for some cycle $\gamma$ of $G$ (Proposition 8.3); the cycle $\gamma$ and the wall are also computed. With the notations of the proof of Proposition 8.3, there exists a two-sided circle $C$ of the subdivided circular wall $\dot{W}$ such that the edges of $O$ that have an endpoint of $C$ leave $C$ all on the same of $C$. Moreover, clearly, we can compute such a circle $C$ in linear time. Once this is done, as in the proof of Proposition 8.3, from the proper embedding of a $5c$-homeomorph of $G$, augmented with at most $c$ isolated vertices, we can compute, in linear time, a proper embedding of a $5c$-homeomorph of $G$, augmented with at most $c$ isolated vertices, into $\mathscr{C}$. ∎
@@
We now have a proper embedding of a $5c$-homeomorph of $G^{\prime}$, augmented with at most $2c$ vertices and at most $3c+k$ edges, into $\mathscr{C}$, satisfying the property that this embedding, restricted to the detached surface $\mathscr{S}$, is cellular. If necessary, we subdivide further the edges to obtain a proper embedding of the graph $G^{\prime}$, in which each edge is subdivided at most $5c$ times, and then augmented with at most $2c$ vertices and at most $3c+k$ edges. We dissolve degree-two vertices that are not part of $G$ and not on singular points. There remains, if necessary, to subdivide the edge of $G^{\prime}$ corresponding to the path component of $G$ at most $2k$ times, so that it now contains all the path components and isolated vertices of $G$, in order to obtain the desired embedding; this adds at most $k$ edges that are not part of $G$.
Eventually, we have a graph $H$ obtained from $G$ by augmenting it with at most $2c$ vertices and at most $3c+2k$ edges, and performing at most $c$ edge subdivisions (the corresponding vertices being on singular points of $\mathscr{C}$). Moreover, the images of the vertices of $H$ cover the singular points of $\mathscr{C}$. Also, $H$ corresponds to a cellular embedding on the detached surface $\mathscr{S}$, so it can be encoded by a combinatorial map as claimed, and similarly one can encode the embedding $H$ restricted to the isolated edges of $\mathscr{C}$. ∎

### v3

@@ 8.1 Finding a large planar part
We claim the following: If at least $1000g$ inner vertices of $W_{1}$ are connected to the outside, then $G$ has genus at least $g$. The strategy is similar to the argument in Kociumaka and Pilipczuk [27, Lemma 5.3]; we summarize the proof. If at least $1000g$ inner vertices of $W_{1}$ are connected to the outside, then a set $U$ of $g$ inner vertices of $W_{1}$ are connected to the outside, and at pairwise distance at least 16 in $W_{1}$. This implies that $G$ contains, as a minor, the graph $J_{g}$ obtained from $g$ copies of $K_{5}$ by subdividing an edge from each copy with a degree-two vertex and identifying these $g$ new vertices into a single vertex, the apex of $J_{g}$; see Figure 8.2. This graph has genus at least $g$ [34, Theorem 1]. This proves the claim.
@@
Figure 8.2: Illustration of the proof of the claim in Proposition 8.1. The wall $W_{1}$ (left), and a schematic view of the $J_{p}$ minor (right), illustrated with $g=6$. Let $K_{5}^{-}$ be the graph $K_{5}$ with one edge removed. If $W_{1}$ has many vertices connected to the outside, then a set $U$ of at least $g$ vertices connected to the outside (represented as big disks on the left) are pairwise distant. In the neighborhood of these vertices, we build a $K_{5}^{-}$ minor of $\dot{W}_{1}$ for each vertex in $U$, in which the vertex of $U$ is the “central” vertex (the thick paths need to be contracted to obtain copies of $K_{5}^{-}$). Because the vertices of $U$ are pairwise distant, these minors are pairwise disjoint, and moreover the graph $X$ consisting of the subdivided wall $\dot{W}$ minus the union of these $K_{5}^{-}$ minors is connected. From these $K_{5}^{-}$ minors, we build a $J_{g}$ minor of $G$ (right) as follows. The apex results from the contraction of $X$. Moreover, recall that each vertex of $U$ is connected to the outside; each light thick edge, connecting some vertex $u\in U$ to $X$, is obtained by contracting the corresponding path from $u$ to a vertex of $X$, except its first edge.
@@ 8.2 Finding an irrelevant vertex
Let $\mathscr{C}$ be a 2-complex with $c\geq 1$ simplices. Let $G$ be a graph and $\gamma$ be a cycle in $G$ such that one connected component of $G-\gamma$ is planar and contains a subdivision of the $(200c\times 200c)$-wall. Let $v$ be the central vertex of this wall. Then $G$ is embeddable on $\mathscr{C}$ if and only if GvG-v is.
The proof is inspired from an article by Kociumaka and Pilipczuk [27, Section 5.3], but we provide a slightly different argument, also because we need to handle a 2-complex, not a surface. We use the topological concept of (free) homotopy, see, e.g., Stillwell [47]; for completeness, we recall the definition. A closed curve on a surface $\mathscr{S}$ is a continuous map from the unit circle $S^{1}$ into $\mathscr{S}$. Two closed curves $c_{0}$ and $c_{1}$ are homotopic if, intuitively, one can be deformed to the other or to its reversal; formally, this means that there is a continuous map $h:[0,1]\times S^{1}\to\mathscr{S}$ such that for each $t\in[0,1]$, we have $h(0,t)=c_{0}(t)$, and either for each $t$, $h(1,t)=c_{1}(t)$ or for each $t$, $h(1,t)=c_{1}(1-t)$.
@@
On a surface of genus $g$, a set of disjoint, simple closed curves belong to at most $3g+2$ (free) homotopy classes (here we regard a closed curve and its reversal to be in the same homotopy class).
@@
We need some preparations. A circular wall $W$ [27, Figure 8] of height $h$ and circumference $\ell$ is a 3-regular graph that, in some embedding of $W$ in the plane, is represented as the union of $h$ vertex-disjoint cycles, called circles, organized in a concentric way, such that any two consecutive circles are connected by $\ell$ radial edges; in any two successive layers, the radial edges are interleaved.
Let $H$ be the connected component of $G-\gamma$ that is planar and contains a subdivision of the $(200c\times 200c)$-wall. By 3-connectivity, this wall has a unique combinatorial embedding in the plane, up to symmetry and up to the choice of the outer (infinite) face; we denote by $\varphi$ a plane embedding of $H$ such that the restriction of $\varphi$ to the wall is the natural embedding, in which the outer face of the wall has the largest degree. In particular, $v$ is the central vertex of this wall.
@@
Then $H$ contains a subdivision $\dot{W}$ of a circular wall $W$ of height $11c+4$ and circumference $2c+1$, so that $v$ is located inside the inner circle of this circular wall in $\varphi$; see Kociumaka and Pilipczuk [27, Figures 8 and 9].
@@
If $G$ is embeddable on $\mathscr{C}$, then obviously GvG-v is also embeddable on $\mathscr{C}$; the hard part is the reverse direction. So let us consider an embedding $\psi$ of GvG-v on $\mathscr{C}$. This induces embeddings of $\dot{W}$ and $W$ on $\mathscr{C}$. At most $c$ vertices or edges of $W$ are mapped, under $\psi$, to a singular vertex of $\mathscr{C}$, and the isolated edges of $\mathscr{C}$ intersect the images of at most $c$ edges of $W$. Thus, we obtain a subdivision $\dot{W}^{\prime}$ of a circular wall $W^{\prime}$ of height $9c+4$ and circumference one that does not intersect any singular vertex or isolated edge of $\mathscr{C}$.
Let $C_{1},\ldots,C_{9c+4}$ be the cycles in GvG-v that correspond to the circles of $W^{\prime}$, such that in $\varphi$, they are nested in this order, $C_{1}$ being the innermost cycle. Let $C$ be the union of these cycles. A bridge (of $C$) is either (1) a connected component of GCG-C, together with the edges joining it to $C$, or (2) an edge of GCG-C whose endpoints are on $C$. For a bridge of $C$, the vertices of $C$ that are endpoints of edges in the bridge are the attachment points of the bridge. For each integer $i$, let $B_{i}$ be the set of bridges all of whose attachment points belong to $C_{i}$; in particular, $v$ belongs to a bridge in $B_{1}$. Moreover, for each integer $i$, let $B_{i+\nicefrac{{1}}{{2}}}$ be the set of bridges having at least one attachment point on $C_{i}$ and at least one on $C_{i+1}$. The existence of $\varphi$ implies that there are no other bridges.
We claim that, for some integer $i_{1}$, the cycle $C_{i_{1}}$ is two-sided and the edges of the bridges in $B_{i_{1}+\nicefrac{{1}}{{2}}}$ leave $C_{i_{1}}$, in $\psi$, all on the same side of $C_{i_{1}}$. We prove the claim in the next few paragraphs.
@@
Let $\mathscr{S}$ be the detached surface of $\mathscr{C}$, and let $\hat{\mathscr{S}}$ be the surface obtained by attaching a handle to every boundary component of $\mathscr{S}$ and to a neighborhood of each point of $\mathscr{S}$ not on its boundary that corresponds, in $\mathscr{C}$, to a singular vertex. At most $2c$ handles were attached, so $\hat{\mathscr{S}}$ has a natural cellular graph embedding with at most $3c$ edges, and thus has genus at most $3c$. Moreover, $\psi$ naturally corresponds to an embedding of $\dot{W}^{\prime}$ and $W^{\prime}$ on $\hat{\mathscr{S}}$, which we still denote by $\psi$.
By Lemma 8.4, since $\hat{\mathscr{S}}$ has genus at most $3c$, and $W^{\prime}$ has height (33c+2)+2(3*3c+2)+2, there are two cycles $C_{i_{0}}$ and $C_{i_{1}}$, with $i_{0}<i_{1}<i_{2}:=(3*3c+2)+2$, that are homotopic in $\hat{\mathscr{S}}$ under $\psi$. They are two-sided, because any two one-sided homotopic closed curves cross. To prove our claim, we distinguish two cases, implicitly using the fact that, as testified by $\varphi$, each bridge in $B_{i_{1}+\nicefrac{{1}}{{2}}}$ is connected to $C_{i_{2}}$ in $G-(C_{i_{0}}\cup C_{i_{1}})$:
  • Assume that $\psi(C_{i_{1}})$ is contractible on $\hat{\mathscr{S}}$; it thus bounds a disk on $\hat{\mathscr{S}}$, and thus separates $\mathscr{C}$ into two pieces. Thus, the edges of the bridges in $B_{i_{1}+\nicefrac{{1}}{{2}}}$ leave $C_{i_{1}}$, in $\psi$, on the side of $C_{i_{1}}$ corresponding to the piece that contains $C_{i_{2}}$, as desired.
@@
  • Otherwise, the closed curves $\psi(C_{i_{0}})$ and $\psi(C_{i_{1}})$ are homotopic on $\hat{\mathscr{S}}$ but non-contractible, and thus bound an annulus in $\hat{\mathscr{S}}$ [20, Lemma 2.4], and thus also in $\mathscr{C}$; we denote by $A$ the annulus on $\mathscr{C}$. There are two subcases:
@@
    • Assume that, under $\psi$, the annulus $A$ does not contain $C_{i_{2}}$. Then, $B_{i_{1}+\nicefrac{{1}}{{2}}}$ cannot enter $A$, which implies the claim;
    • otherwise, under $\psi$, $B_{i_{1}+\nicefrac{{1}}{{2}}}$ must lie entirely inside $A$, which also implies the claim.
So the claim is proved.
To conclude the proof, we first consider the restriction $\psi^{\prime}$ of $\psi$ to $G\setminus(C_{\leq i_{1}-1}\cup B_{\leq i_{1}})$. (For each integer $i$, $C_{\leq i}$ denotes the union of the cycles $C_{j}$, $1\leq j\leq i$; similarly, for each integer or half-integer $i$, $B_{\leq i}$ denotes the union of the bridges $B_{j}$, $j$ being an integer or a half-integer between 1 and $i$.)
Next, we extend $\psi^{\prime}$ to an embedding $\psi^{\prime\prime}$ of $G\setminus B_{i_{1}}$ into $\mathscr{C}$, by mapping $C_{\leq i_{1}-1}\cup B_{\leq i_{1}-\nicefrac{{1}}{{2}}}$ close to $C_{i_{1}}$, on the side of $C_{i_{1}}$ that is not already used, as in $\varphi$.
Finally, there remains to extend $\psi^{\prime\prime}$ to $B_{i}$. For this, we note that, under $\varphi$, the set of edges of $G$ incident to $C_{i_{1}}$ is divided into two parts according to the side of $C_{i_{1}}$ they leave $C_{i_{1}}$, and that, by construction, this partition is the same under $\psi^{\prime\prime}$. We can thus augment $\psi^{\prime\prime}$ to an embedding of all of $G$ into $\mathscr{C}$, by embedding $B_{i_{1}}$ close to $C_{i_{1}}$ as it is done in $\varphi$. ∎
@@
    a rooted branch decomposition of $G$ of width $g^{O(1)}$;
@@
In the first two cases, we are done. In the third case, by applying Proposition 8.3, we obtain a vertex $v$ such that $G$ embeds on $\mathscr{C}$ if and only if GvG-v does. By iterating the same procedure a number of times that is at most the number of vertices of $G$, we necessarily reach case (1) or (2), which concludes. ∎
@@ 9 Computing an embedding
  • the images of the vertices of $H$ cover the singular vertices of $\mathscr{C}$;
  • the restriction of $H$ to the detached surface $\mathscr{S}$ is specified by (1) a cellular embedding of a graph $H^{\prime}$, represented by its (standard) combinatorial map, such that $E(H^{\prime})\subseteq E(H)$, (2) a map from $V(H^{\prime})$ to $V(H)$ describing how the vertices of $H^{\prime}$ are identified when attaching the points of $\mathscr{S}$ to recover the singular vertices of $\mathscr{C}$, and (3) a map from the images of the singular vertices of $\mathscr{C}$ on $\mathscr{S}$ to $V(H^{\prime})$, to specify which vertex of $H^{\prime}$ occupies that singular vertex. Moreover, we have $|V(H^{\prime})|\leq|V(H)|+c$;
  • the image of each isolated edge $e$ of $\mathscr{C}$ is the disjoint union of the images of some vertices and edges of $H$, and of some connected parts of $e$ not in the image of $H$; we can represent these objects by listing them in order along $e$.
We remark that it is not a severe restriction to assume that $\mathscr{C}$ has no 3-book, since any graph embeds on a 3-book (Proposition 3.1), and no isolated vertex, since only isolated vertices of $G$ can be embedded on such isolated vertices. Also, we remark that the graph $H$ is needed because we may need to add edges to make the embedding cellular on the detached surface, and we may need to subdivide edges in order to avoid any edge to go through a singular vertex.
@@
    Since $G$ embeds on $\mathscr{C}$, for some complex $\mathscr{C}_{i}$, we have computed a sparse bounding graph for $G^{\prime}$ on $\mathscr{C}_{i}$ at the root of the rooted branch decomposition. We can use backtracking: at each node $\alpha$ of the rooted branch decomposition, we store the embedding of the union of $G_{\alpha}$ and of its bounding graph, which is a proper embedding. Eventually, the embedding of $G^{\prime}$ at the root of the recursion is not necessarily proper, but it becomes proper if we add isolated vertices mapped to singular vertices that are not covered by the embedding of $G^{\prime}$. Thus, from $(G^{\prime},\mathscr{C}_{i})$, we can return a proper embedding of the graph $G^{\prime}$, augmented with at most $c$ isolated vertices, into $\mathscr{C}_{i}$.
@@
In Theorem 1.1, if $\mathscr{C}$ has no 3-book and no isolated vertex, $G$ has no isolated vertex and at most one connected component that is a path, and $G$ embeds into $\mathscr{C}$, then one can compute a proper embedding of a $5c$-homeomorph of $G$, augmented with at most $c$ isolated vertices, into $\mathscr{C}$ (using the data structure of Section 3.4), without overhead in the asymptotic running time.
@@
The decision algorithm iteratively removes vertices to $G$, getting a smaller equivalent instance. For the algorithm that actually computes an embedding, we now need recursion. The base case, for bounded branchwidth, is given by Lemma 9.2.
In the general case, we have computed a cycle $\gamma$ of $G$ such that one connected component of $G-\gamma$ is planar and contains a subdivision of the $(200c\times 200c)$-wall; this wall, and the center $v$ of this wall, have also been computed (see Proposition 8.3). We assume that we have a proper embedding $\psi$ of a $5c$-homeomorph of GvG-v, augmented with at most $c$ isolated vertices, into $\mathscr{C}$. We need to recover a proper embedding of a $5c$-homeomorph of $G$, augmented with at most $c$ isolated vertices, into $\mathscr{C}$.
The algorithm follows the proof and notation of Proposition 8.3. Starting from the above wall, we compute the subdivision $\dot{W}$ of the corresponding circular wall $W$. From the embedding $\psi$, we obtain a subdivision $\dot{W}^{\prime}$ of a circular wall $W^{\prime}$ of height $9c+4$ and circumference one that does not intersect any singular vertex or isolated edge of $\mathscr{C}$. In linear time, we compute $\varphi$, and also the bridges of $C$ and their partition into the sets $B_{i}$.
In the proof of Proposition 8.3, we proved that, for some $i_{1}$, the cycle $C_{i_{1}}$ is two-sided and the edges of the bridges in $B_{i_{1}+\nicefrac{{1}}{{2}}}$ leave $C_{i_{1}}$, in $\psi$, all on the same side of $C_{i_{1}}$. From the $B_{i}$s, we can easily compute such a value of $i_{1}$ in linear time. Once this is done, as in the proof of Proposition 8.3, from the embeddings $\varphi$ and $\psi$ we can compute a proper embedding of a $5c$-homeomorph of $G$, augmented with at most $c$ isolated vertices, into $\mathscr{C}$; this takes linear time. ∎
@@
We now have a proper embedding of a $5c$-homeomorph of $G^{\prime}$, augmented with at most $2c$ vertices and at most $3c+k$ edges, into $\mathscr{C}$, satisfying the property that this embedding, restricted to the detached surface $\mathscr{S}$, is cellular. If necessary, we subdivide further the edges to obtain a proper embedding of the graph $G^{\prime}$, in which each edge is subdivided at most $5c$ times, and then augmented with at most $2c$ vertices and at most $3c+k$ edges. We dissolve degree-two vertices that are not part of $G$ and not on singular vertices. There remains, if necessary, to subdivide the edge of $G^{\prime}$ corresponding to the path component of $G$ at most $2k$ times, so that it now contains all the path components and isolated vertices of $G$, in order to obtain the desired embedding; this adds at most $k$ edges that are not part of $G$.
Eventually, we have a graph $H$ obtained from $G$ by augmenting it with at most $2c$ vertices and at most $3c+2k$ edges, and performing at most $c$ edge subdivisions (the corresponding vertices being on singular vertices of $\mathscr{C}$). Moreover, the images of the vertices of $H$ cover the singular vertices of $\mathscr{C}$. Also, $H$ corresponds to a cellular embedding on the detached surface $\mathscr{S}$, so it can be encoded by a combinatorial map as claimed, and similarly one can encode the embedding $H$ restricted to the isolated edges of $\mathscr{C}$. ∎